Solve the literal equation for the specified variable.
P=2L+2W for W
A)W=d-2L
B)W=P-2L/2
C)W=P-L
D)W=P-L/2

Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

Solve the literal equation for the specified variable.
P=2L+2W for W
-----
2W = P-2L
----
W = (P-2L)/2
Cheers,
Stan H.
--------------------
A)W=d-2L
B)W= (P-2L)/2
C)W=P-L
D)W=P-L/2

Answer by mananth(16946)


P=2L+2W for W
P-2L=2W
(P-2L)/2=w